40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2668     Entry time: Thu Mar 11 17:51:04 2010
Author: Koji 
Type: Update 
Category: SUS 
Subject: Recent status of SOSs 

Jenne, Koji

Recent status of SOSs:

We completed one of the suspension (ITMY).

ITMX: 6 Magnets, standoffs, and guide rod glued / balance to be confirmed / needs to be baked
ITMY: 6 Magnets, standoffs, and guide rod glued / balance confirmed / needs to be baked
SRM: 6 Magnets, one standoff, and guide rod glued,  / waiting for the release from the gluing fixture.
PRM: one standoff, and guide rod glued / waiting for the magnet gluing. 

We think we solved all the problems for hanging the suspensions.

--- Magnet gluing fixture ---

  • There is the two kinds of fixtures. Neither does work propery in the original form!
  • The height of the side magnets should be finely adjusted by changing the teflon sheets beneath the optics in the fixture.
  • Be aware of the polarity of the fixture in terms of the side magnets
  • Wrongly glued magnets (and others) can be removed by a razor blade with some amount of acetone.
  • The pickle picker frequently knocks the magnets down during the release. Don't s be down in the dumps too much.

--- Suspending the mirror ---

  • The wire winches must be carefully attached to the suspension tower such that the wires are not streached during fastening the clamps.
  • There are a couple variations of the drawings for SOS. The one we have has #4-40 for the earthquake stops at the bottom.
    Zach and Mott made the EQ stops with the right size.
ELOG V3.1.3-